OpenStudy (pphalke):

If a right triangle has one leg that measures 12 cm, and a hypotenuse that measures 13 cm, how long is the other side?

Mehek (mehek14):

Use the Pythagorean Theorem \[a^2+b^2=c^2 \] One leg is 12 so put 12 for a and 13 is the hypotenuse so put 13 for c \[12^2+b^2=13^2\] What is \[12^2=?\\13^2=?\]

OpenStudy (pphalke):

12^2 is 144 and 13^2 is 169

Mehek (mehek14):

Correct \[144+b^2=169\] We have to isolate b. Subtract \[169-144\]

OpenStudy (pphalke):

25 and sq rt of 25 is 5

Mehek (mehek14):

Yes so the length is 5